"Part of the income generated by an ice cream machine is saved to purchase a new machine in the future. If $5,769 is deposited annually at 4% interest, how many years must the current machine last before a new machine costing $60,866 can be purchased? The first deposit will occur exactly one year from now. (Enter your answer in number of years as an integer.)"
Answer #1

ANSWER:

I = 4%

N = ?

FV = $60,866

ANNUAL DEPOSIT = $5,769

FV = ANNUAL DEPOSIT(F/A,I,N)

60,866 = 5,769(F/A,4%,N)

60,866 / 5,769 = (F/A,4%,N)

10.55 = (F/A,4%,N)

Looking into compounding factor tables we get that n is between 8 and 9 years and solving further we get that n is 8.97 years.

so it will take 9 years (rounded off) before the current machine expires.